# Efficient onboard multi-task AI architecture based on self-supervised learning

### v1

@@ I Introduction
An emerging paradigm [1] is to move part of the processing directly onboard the satellite, so that it can detect potentially critical situations in real time and trigger early warnings whose quick transmission to the ground segment is prioritized. This process requires an at least partial onboard formation of image products, followed by image analysis; a full onboard pipeline of optical and SAR image formation, analysis and alert generation has been demonstrated in [2]. Achieving this goal requires facing a challenging tradeoff between the quality of the detection, its speed, and the computational constraints of onboard platforms dictated by the strict power budgets of satellites. Moreover, multispectral imagers are capable of detecting several phenomena of interest, such as floods, fires, clouds, marine debris, and many more, leading to the capability of addressing multiple tasks at the same time. Indeed, onboard multitask inference would not only allow to monitor multiple critical phenomena at the same time but also synergistically improve the entire platform. For example, segmentation of clouds could be used to optimize the onboard compression algorithm by lowering the data rate for cloud-covered areas.
@@
Targeting a Sentinel2-like multispectral imager, we present a lightweight backbone design, trained with a self-supervised procedure that promotes features with fine spatial resolution, and three parallel heads for as many sample tasks of interest: cloud segmentation, flood detection, and marine debris classification. We show that our lightweight and modular approach to mission design provides inference accuracy close to that of high-complexity state-of-the-art models and higher throughput. We also report experimental results on a 7W embedded system, showing an excellent tradeoff between quality, throughput and power consumption.
@@ II-A Self-Supervised Learning (SSL)
Self-Supervised Learning (SSL) has emerged in the last years as a poweful paradigm in deep learning, aiming at learning good representations that capture intrinsic data features without relying on human-labeled annotations. This is critical in remote sensing due to the constant scarcity of labeled data and the abundance of unlabeled imagery. Contrastive learning is currently one of the most successful SSL techniques where informative representations emerge from minimizing the distance between the feature-space embedding of the same image subjected to two distinct random augmentations (positive pair) while maximizing the distance between representations derived from distinct images (negative pairs). Works such as SimCLR [3] highlight the efficacy of contrastive methods in learning resilient and generalizable representations, although they are not without flaws. In fact, they often need a large batch size to work properly and also the handling of negative pairs needs to be carefully managed. The need of a large batch size was partially solved by MoCo (Momentum Contrast) [4], using a momentum moving average encoder. Newer methods like BYOL [5] also address the problem of creating truly negative pairs by relying only on positive ones. Furthermore, He et al. [6] observed that image-level learning does not always provide good representations for sub-pixel level tasks such as semantic segmentation and object detection limiting the effectiveness of SSL only to image-level tasks such as classification. For this reason, research has started investigating “Dense SSL” techniques [7, 8, 9] capable of learning more fine-grained features.
@@ III-A Mission Vision
The approach towards the design of neural network components described in later sections stems from ideas about the specific goals and planning requirements of a hypothetical mission. We envision a Sentinel2-like multispectral imager with additional capabilities provided by onboard AI.
In particular, the first novel capability would be AI-assisted onboard compression. It is known that image compression methods, including existing standards for hyper- and multispectral images, can be aided by cloud detection [21] to provide pixel-level maps of regions where compression quality can be lowered to significantly save data rate. Cloud segmentation is therefore a desirable task to be included for any onboard AI capability. Furthermore, the second capability of interest is the generation of alerts to be delivered to the ground segment with low latency when specific phenomena, such as natural disasters, are detected. For this capability, it is desirable to produce both pixel-level segmentation maps (e.g., to detect the extent of flooded areas) as well as whole-image classification labels (e.g., to detect debris presence and its type, or presence of active fires). These requirements clearly outline the need to have features with fine spatial granularity so that the segmentation tasks can be solved effectively.
@@ III-B Self-supervised Backbone
Concerning the architecture design, we suppose that four spectral bands (RGB and NIR, corresponding to bands 2,3,4,8 of Sentinel-2) are used as input. The choice of bands is a tradeoff between the tasks to be solved and computational complexity. Using all the available spectral bands could provide a more flexible backbone, but also increase computational complexity. For the sample tasks explored in this paper, the RGB and NIR bands provide adequate information, but other tasks could require additional bands. For instance, SWIR could be useful to target fire detection and one could imagine to implement a slightly more complex head that takes as input the backbone features as well as the pixels of a new (e.g. SWIR) channel, and combines them to produce output for this specific task.
@@ III-C Task-specific Heads
The general framework outlined in this paper enables a large variety of applications to be addressed thanks to the features extracted from the backbone. In order to evaluate the effectiveness of our design, we tested three tasks which can be relevant for onboard inference: cloud cover segmentation [32], floods segmentation [33], and marine littering multi-label classification [34]. The following datasets have been used for the sample tasks.
@@
  • MARIDA [34]: Marine Debris Archive (MARIDA) is a dataset for the detection of marine debris. The dataset includes 1381 Sentinel-2 multispectral images of 256x256 pixels, which distinguishes marine debris from various coexisting marine features, including Sargassum macroalgae, ships, natural organic material, waves, wakes, foam, different water types (e.g., clear water, turbid water, sediment-laden water, shallow water), and clouds. We use this dataset for whole-image multi-label classification where the entire input image is classified in one of 11 classes (different water types are aggregated into one class, reducing the number from the original 15 to 11).
Each dataset underwent a cleaning procedure, systematically removing all images containing exclusively unlabeled or no-data pixels.
@@ IV Experimental results
In this section, we initially validate the proposed design against state-of-the-art methods comparing their complexity as well as inference performance (in both FP32 and INT8 precision) on the three sample tasks presented in Sec. III-C. Then, we use low-power hardware to analyze how the propose design fares in terms of total energy consumption and throughput. Lastly, we present ablations by comparing a set of experiments featuring backbones of varying sizes.
@@ IV-A Implementation details
The SSL training of the backbone spanned 550 epochs, employing a learning rate of $3\times 10^{-4}$ and a batch size of 3400 parallelized over 4 Nvidia A100 GPUs. After the completion of SSL training, we extracted the GhostNetV2 from the online branch and employed it as backbone feature extractor. Following a procedure similar to that presented in [7], we generated two distinct augmented views of a single image: $I_{s}$ resulted solely from spatial transformations encompassing horizontal and vertical flips, random rotations (90°, -90°, 180°, -180°), and random cropping. While, $I_{c}$ was derived by applying a combination of spatial and spectral transformations, including color jittering, intensity manipulation, gaussian blurring, and solarization. We determined that hyperparameter $\lambda=0.1$ offered the optimal balance between the global and local contrastive losses. This choice resulted in a balanced performance across sub-pixel level and image-level tasks. Unless otherwise stated, the GhostNetV2 architecture uses multiplier $\alpha=1.6$ in the choice of number of features. Finally, we specify that in order to perform network quantization in a fair way, the Intel Neural Compressor [38] library was used for all models. Each model was quantized in the same way, using a static configuration for post-training quantization, with arithmetic entirely on 8-bit integers and calibrating each model with a calibration sample of 100 elements.
@@ IV-B Model Comparison and Performance Evaluation
Architecture mIOU (FP32) mIOU (INT8)
GhostNetv2 + Segmentation Head (SSL) 82.33 81.7
GhostNetv2 + Segmentation Head (SL) 83.75 83.41
DeepLabV3 + MobileNetV3-L 83.47 81.8
HRNet18 84.67 84.57
UNet + MobileViT-S 83.55 50.74
The tasks outlined in Sec. III-C include two semantic segmentation tasks (cloud cover segmentation [32], flood segmentation [33]) and one multi-label classification task (marine litter [34]). Accordingly, we select some baselines and state-of-the-art models for segmentation and classification as terms of comparison. In particular, concerning segmentation we select DeepLabV3 [39] with MobileNetV3-L [28] backbone for a well-known and efficient baseline, HRNet18 [30] as the high-complexity state-of-the-art model, and UNet [40] with MobileVit-S as backbone [28] as a recent approach leveraging the representational power of Transformers, albeit with an eye to complexity. Concerning classification, we consider ResNet50 [41] as baseline, MobileNetV3 [28] as lightweight method and ViT B-16 [42] as high-complexity state-of-the-art model.
A summary of the computational complexity of various methods is reported in Table I. For ease of comparison we also include the proposed design with a single head in addition to the multitask case. We can see that the proposed design is generally lighter while having a similar number of parameters (except for the large classification models). As a reference, our single segmentation head has 401K parameters and requires 92K MACs, 187M FLOPs for inference on a $512\times 512\times 4$ image provided to the backbone, while a single classification head has 743K parameters and requires 742K MACs, 1.88M FLOPs for inference on a $512\times 512\times 4$ image provided to the backbone. These are comparatively small with respect to the backbone.
@@
Architecture B. mIOU (FP32) B. mIOU (INT8) M. mIOU (FP32) M. mIOU (INT8) mF1 (FP32) mF1 (INT8)
GhostNetv2 + Segmentation Head (SSL) 40.32 39.78 65.41 65.13 50.45 49.66
GhostNetv2 + Segmentation Head (SL) 42.95 42.76 67.15 67.03 53.71 53.26
DeepLabV3 + MobileNetV3-L 41.03 34.32 66.27 62.78 51.39 43.62
HRNet18 54.68 54.62 74.00 74.00 65.72 65.51
UNet + MobileViT-S 59.65 14.75 76.83 47.60 70.25 19.18
@@
Concerning training, our proposed backbone underwent SSL training on the BigEarthNet dataset and then its weights were frozen to the pretrained values. Heads were trained on a supervised way on each task dataset, without finetuning the backbone. In order to provide a fair comparison, the other methods were pretrained on ImageNet [43] and finetuned on the task datasets. We remark that freezing the backbone as dictated by our design goals is nevertheless penalizing compared to full finetuning. For this reason and comparison sake, we also report a benchmark in which the proposed model is fully finetuned for a specific task (this will be marked in the following as “SL” - supervised learning, in contrast to the “SSL” configuration for the frozen backbone), after the SSL pretraining.
Tables II, III, IV report performance comparisons for the cloud cover segmentation, flood segmentation and marine litter classification, respectively.
@@
For what concerns cloud cover segmentation, we ensure consistency with published results [44] by adopting the mean Binary Intersection-over-Union (mIOU) metric. It is noteworthy that all models exhibit relatively similar performance, within about 2 percentage points of variation in mIOU between the best and worst, both in FP32 and INT8 quantization, with the exception of UNet + MobileViT-S which suffers greatly from quantization. Furthermore, it is interesting to notice that the SL version of our architecture is only marginally better than the SSL version, suggesting that the design constraint of freezing the backbone may not have a big impact.
@@
For what concerns the floods segmentation task, we decided to evaluate three different quality metrics: Binary Intersection-over-Union, Multiclass Intersection-over-Union (with 2 classes, flood-water and no-water) and the F1-Score. The reasoning behind including two different mIOU metrics lies in a more comprehensive evaluation, since the binary IoU gives us an assessment of how well the “true” class (i.e., floods) was predicted, the multiclass IoU for two classes offers a nuanced view of the performance of individual classes within the segmentation task. The results on this task show that the higher complexity models are generally superior, while the proposed architecture provides better performance than the direct low-complexity alternative (DeepLabV3 + MobileNetV3-L). We attribute this to the relatively small number of parameters of the proposed architecture. Indeed, scaling experiments reported in Table VII suggest that a larger model would improve performance in exchange of speed.
@@
Lastly, the marine littering task addressed a problem of multilabel classification instead of segmentation, providing an analysis of how well the proposed model can address heterogeneous tasks. In this task, we notice that the proposed method is very close to the best FP32 results, and it is the best overall in INT8. It is worth remarking that the MARIDA dataset is very small (only 1300 images of size $256\times 256\times 4$ to cover 11 different classes), leading some highly-complex methods to overfit when finetuned.
@@
Model Tasks Inference Latency (FP32) Inference Latency (INT8) Avg Power
GhostNetv2 + 3 parallel heads (Ours) SSC 56.77 ms 34.67 ms 6000 mW
DeepLabV3 + MobileNetV3-L (DLMN) S 39.52 ms 15.70 ms 5800 mW
HRNet18 (HR) S 118.07 ms 47.72 ms 6300 mW
UNet + MobileViT Small (mViT) S 117.79 ms 82.94 ms 6500 mW
MobileNetV3 Large (MN) C 18.61 ms 9.19 ms 5800 mW
ResNet50 (RN50) C 45.98 ms 15.00 ms 6400 mW
ViT B-16 (ViT) C 364.79 ms 296.27 ms 6600 mW

### v2

@@ I Introduction
An emerging paradigm [1] is to move part of the processing directly onboard the satellite, so that it can detect potentially critical situations in real time and trigger early warnings whose quick transmission to the ground segment is prioritized. This process requires an at least partial onboard formation of image products, followed by image analysis; a full onboard pipeline of optical and SAR image formation, analysis and alert generation has been demonstrated in [2]. Achieving this goal requires facing a challenging tradeoff between the quality of the detection, its latency, and the computational constraints of onboard platforms dictated by the strict power budgets of satellites. Moreover, multispectral images are capable of detecting several phenomena of interest, such as floods, fires, clouds, marine debris, and many more, leading to the capability of addressing multiple tasks at the same time. Indeed, onboard multitask inference would not only allow to monitor multiple critical phenomena at the same time but also synergistically improve the entire platform. For example, segmentation of clouds could be used to optimize the onboard compression algorithm by lowering the data rate for cloud-covered areas.
@@
To summarize, this paper presents the following key novel contributions:
  • we study how to design the onboard AI system for an Earth observation mission required to address multiple tasks, analyzing the entire framework needed to accomplish this goal, the technical details of the individual components, and presenting novel methodologies;
  • we show that a desirable design pretrains a backbone neural network with a self-supervised strategy, and, in contrast with classical self-supervised learning (SSL) literature, keeps it frozen to allow processing multiple tasks in parallel;
  • we propose a novel self-supervised learning method that can extract features with both fine and coarse spatial resolution, outperforming existing SSL methods in remote sensing;
  • we evaluate the effects of quantization on SSL pretrained models, a topic rarely explored in the SSL literature;
  • we demonstrate a lightweight and modular design that provides inference accuracy close to that of high-complexity state-of-the-art models while achieving higher throughput on three tasks for onboard computing (clouds segmentation, floods segmentation and marine debris classification)
  • we demonstrate that the proposed design has an excellent trade-off between quality, throughput, and power consumption on a low-power 7W embedded system.
@@ II-A Self-Supervised Learning (SSL)
Self-Supervised Learning (SSL) has emerged in the last years as a poweful paradigm in deep learning, aiming at learning good representations that capture intrinsic data features without relying on human-labeled annotations. This is critical in remote sensing due to scarcity of labeled data and the abundance of unlabeled imagery. Contrastive learning is currently one of the most successful SSL techniques where informative representations emerge from minimizing the distance between the feature-space embedding of the same image subjected to two distinct random augmentations (positive pair) while maximizing the distance between representations derived from distinct images (negative pairs). Works such as SimCLR [3] highlight the efficacy of contrastive methods in learning resilient and generalizable representations, although they are not without flaws. In fact, they often need a large batch size to work properly and also the handling of negative pairs needs to be carefully managed. The need of a large batch size was partially solved by MoCo (Momentum Contrast) [4], using a momentum moving average encoder. Newer methods like BYOL [5] also address the problem of creating truly negative pairs by relying only on positive ones. Furthermore, He et al. [6] observed that image-level learning does not always provide good representations for sub-pixel level tasks such as semantic segmentation and object detection limiting the effectiveness of SSL only to image-level tasks such as classification. For this reason, research has started investigating “Dense SSL” techniques [7, 8, 9] capable of learning more fine-grained features.
@@ II-B Efficient Inference with Neural Networks
II-C Onboard AI-Based Processing
In recent years, the advancement of neural networks and AI has extended to onboard satellite processing systems and edge-devices in general, enabling real-time data analysis directly in space. This advancement is particularly significant in the context of remote sensing, where the ability to process large volumes of onboard imagery can dramatically reduce latency, optimize bandwidth utilization, and enable more responsive and autonomous satellite operations. However, onboard AI is faced with several challenges in the design of lightweight and power-efficient systems. Several studies[21] have started demonstrated the feasibility and effectiveness of AI-based onboard processing for remote sensing tasks. Yao et al. [22] was one of the first works to address the challenge of running deep learning models directly onboard satellites, proposing a simple framework for ship detection on small satellites. Notably, Giuffrida et al. [23] demonstrated the integration of AI for onboard data processing in real Earth Observation missions ($\Phi$-Sat-1 by the European Space Agency), showcasing the feasibility of running deep convolutional neural networks on the Intel Movidius Myriad 2 hardware accelerator for real-time cloud detection on hyperspectral images. Ziaja et al. [24] proposed and extensive benchmark of various deep learning models on edge devices for onboard space applications. Růžička et al. [25] introduced a lightweight model for change detection onboard satellites based on Variational Auto-Encoders.
In this paper, as a demonstration of the performance of the proposed design in a low-power setting, we test using an Nvidia Jetson Orin Nano, a Commercial-Off-The-Shelf (COTS) hardware platform. This should be considered as a low-power demonstrator, and not necessarily representative of a real flight implementation. Indeed, we recognize that different space missions may choose different approaches to the integration of deep learning in the onboard computing platforms depending on specific mission characteristics. For instance, they may rely solely on FPGAs, or a combination of FPGAs with GPUs/CPUs or even just a COTS System-on-Chip.
@@ III-A Mission Vision
The approach towards the design of neural network components described in later sections stems from ideas about the specific goals and planning requirements of a hypothetical mission. We envision a Sentinel2-like multispectral imager with additional capabilities provided by onboard AI. In particular, the first novel capability would be AI-assisted onboard compression. It is known that image compression methods, including existing standards for hyper- and multispectral images, can be aided by cloud detection [26] to provide pixel-level maps of regions where compression quality can be lowered to significantly save data rate. Cloud segmentation is therefore a desirable task to be included for any onboard AI capability. Furthermore, the second capability of interest is the generation of alerts to be delivered to the ground segment with low latency when specific phenomena, such as natural disasters, are detected. For this capability, it is desirable to produce both pixel-level segmentation maps (e.g., to detect the extent of flooded areas) as well as whole-image classification labels (e.g., to detect debris presence and its type, or presence of active fires). These requirements clearly outline the need to have features with fine spatial granularity so that the segmentation tasks can be solved effectively.
@@ III-B Self-supervised Backbone
Concerning the architecture design, we suppose that four spectral bands (Red, Green, Blue and Near Infrared) are used as input. The choice of bands is a tradeoff between the tasks to be solved and computational complexity: using all the available spectral bands could provide a more flexible backbone, but also increase computational complexity. For the sample tasks explored in this paper, the RGB and NIR bands provide adequate information, but other tasks could require additional bands. For instance, SWIR could be useful to target fire detection and one could imagine to implement a slightly more complex head that takes as input the backbone features as well as the pixels of a new (e.g. SWIR) channel, and combines them to produce output for this specific task.
@@
Lastly, we want to emphasize that the choice of GhostNetV2 is motivated by being the state-of-the-art model among low-complexity backbones at the time of writing, presenting an excellent tradeoff between complexity and accuracy, which allows us to verify if the onboard multitask AI system can achieve good performance. However, the considerations in this paper are more general and GhostNetV2 could be replaced with any backbone providing multiresolution features, resulting in different tradeoffs between accuracy, latency and power consumption.
@@ III-C Task-specific Heads
The general framework outlined in this paper enables a large variety of applications to be addressed thanks to the features extracted from the backbone. In order to evaluate the effectiveness of our design, we tested three tasks which can be relevant for onboard inference: cloud cover segmentation [37], floods segmentation [38], and marine littering whole-image multi-label classification [39]. The following datasets have been used for the sample tasks.
@@
  • MARIDA [39]: Marine Debris Archive (MARIDA) is a dataset for the classification of marine debris. The dataset includes 1381 Sentinel-2 multispectral images of 256x256 pixels, which distinguishes marine debris from various coexisting marine classes, including Sargassum macroalgae, ships, natural organic material, waves, wakes, foam, different water types (e.g., clear water, turbid water, sediment-laden water, shallow water), and clouds. We use this dataset for whole-image multi-label classification where the entire input image is classified in one of 11 classes (different water types are aggregated into one class, as in the original paper, reducing the number from the original 15 to 11).
@@ IV Experimental results
In this section, we initially validate the proposed design against state-of-the-art methods comparing their complexity as well as inference performance (in both FP32 and INT8 precision) on the three sample tasks presented in Sec. III-C. Then, we use low-power hardware to analyze how the propose design fares in terms of total energy consumption and throughput. Lastly, we present ablations by comparing a set of experiments featuring backbones of varying sizes and comparing how different SSL techniques affect the performance.
@@ IV-A Implementation details
The SSL training of the backbone spanned 550 epochs, employing a learning rate of $3\times 10^{-4}$ and a batch size of 3400 parallelized over 4 Nvidia A100 GPUs. After the completion of SSL training, we extracted the GhostNetV2 from the online branch and employed it as backbone feature extractor. Following a procedure similar to that presented in [7], we generated two distinct augmented views of a single image: $I_{s}$ resulted solely from spatial transformations encompassing horizontal and vertical flips, random rotations (90°, -90°, 180°, -180°), and random cropping. While, $I_{c}$ was derived by applying a combination of spatial and spectral transformations, including color jittering, intensity manipulation, gaussian blurring, and solarization. We determined that hyperparameter $\lambda=0.1$ offered the optimal balance between the global and local contrastive losses. This choice resulted in a balanced performance across sub-pixel level and image-level tasks. Unless otherwise stated, the GhostNetV2 architecture uses multiplier $\alpha=1.6$ in the choice of number of features. Finally, we specify that in order to perform network quantization in a fair way, the Intel Neural Compressor [43] library was used for all models. Each model was quantized in the same way, using a static configuration for post-training quantization, with arithmetic entirely on 8-bit integers and calibrating each model with a calibration sample of 500 elements.
@@ IV-B Model Comparison and Performance Evaluation
Architecture mIoU (FP32) mIoU (INT8) mF1 (FP32) mF1 (INT8)
GhostNetv2 + Segmentation Head (SSL) 82.33 81.7 88.16 87.67
GhostNetv2 + Segmentation Head (SL) 83.75 83.41 89.13 88.85
DeepLabV3 + MobileNetV3-L 83.47 81.8 88.95 87.72
HRNet18 84.67 84.57 89.86 89.79
UNet + MobileViT-S 83.55 50.74 89.04 58.02
The tasks outlined in Sec. III-C include two semantic segmentation tasks (cloud cover segmentation [37], flood segmentation [38]) and one classification task (marine litter [39] multi-label classification). Accordingly, we select some baselines and state-of-the-art models for segmentation and classification as terms of comparison. In particular, concerning segmentation we select DeepLabV3 [44] with MobileNetV3-L [33] backbone for a well-known and efficient baseline, HRNet18 [35] as the high-complexity state-of-the-art model, and UNet [45] with MobileVit-S as backbone [33] as a recent approach leveraging the representational power of Transformers, albeit with an eye to complexity. Concerning classification, we consider ResNet50 [46] as a standard baseline, MobileNetV3 [33] as a lightweight method and ViT B-16 [47] as a high-complexity state-of-the-art model.
A summary of the computational complexity of various methods is presented in Table I. For ease of comparison, we also include the proposed design with a single head alongside the multitask case. The proposed design demonstrates a very low number of FLOPs while maintaining a comparable number of parameters relative to other models, except for the large classification models, which generally require a larger number of parameters. In the context of classification, one might argue that the proposed method has more FLOPs than MobileNetV3, however, when considering the multitask scenario, including also inference for segmentation tasks, the efficient DeepLabV3 + MobileNetV3 framework has more FLOPs than our architecture. Thus, in the multitask setting, our proposed method remains the most efficient. Regarding the complexity of the individual application heads (excluding the backbone): our single segmentation head has 401K parameters and requires 92K MACs, with 187M FLOPs for inference on a $512\times 512\times 4$ image provided to the backbone. Meanwhile, a single classification head has 743K parameters and requires 742K MACs, with 1.88M FLOPs for inference on a $512\times 512\times 4$ image provided to the backbone.
@@
Architecture mIoU (FP32) mIoU (INT8) mF1 (FP32) mF1 (INT8)
GhostNetv2 + Segmentation Head (SSL) 40.32 39.78 50.45 49.66
GhostNetv2 + Segmentation Head (SL) 42.95 42.76 53.71 53.26
DeepLabV3 + MobileNetV3-L 41.03 34.32 51.39 43.62
HRNet18 54.68 54.62 65.72 65.51
UNet + MobileViT-S 59.65 14.75 70.25 19.18
@@
Heads were trained on a supervised way on each task dataset, without finetuning the backbone. In order to provide a fair comparison, the other methods were pretrained on ImageNet [48] and finetuned on the task datasets. We remark that freezing the backbone as dictated by our design goals is nevertheless penalizing compared to full finetuning. For this reason, we also report a benchmark in which the proposed model is fully finetuned for a specific task (this will be marked in the following as “SL” - supervised learning, in contrast to the “SSL” configuration for the frozen backbone), after the SSL pretraining.
Note that all tests conducted in the following section were performed using the original splits provided by the datasets to ensure that the results are comparable with those reported in the datasets’ papers and associated benchmarks. The metrics presented in the tables below were calculated on the test set when available; otherwise, they were calculated on the validation set if the test set was not available in the dataset. The dataset configurations are as follows:
  • Sentinel-2 Cloud Cover Segmentation Dataset: consists of 22,728 total images, with 11,748 in the training set and the remaining 10,980 in the test set.
  • Sen1Floods11: consists of 426 total images, with 256 in the training set, 86 in the validation set, and 89 in the test set.
  • MARIDA: consists of 1,381 images, with 694 in the training set, 328 in the validation set, and 359 in the test set.
Tables II, III, and IV present performance comparisons for cloud cover segmentation, flood segmentation, and marine litter classification, respectively, while in Fig. 4 qualitative results are shown for cloud and flood segmentation tasks, comparing the different segmentation maps obtained from the models with ground truth and the corresponding RGB image. We evaluate the two segmentation tasks using the Binary Intersection-over-Union and the F1-Score, while for the classification task, we compute only the F1-Score. We chose the Binary IoU (hereafter referred to as “mIoU” for brevity) to ensure consistency with published results for cloud segmentation [49] and flood segmentation [38], aligning with the official metric used in both datasets. Additionally, we compute the F1-Score for segmentation to provide a more detailed assessment of the model’s performance across individual classes. In a sensitive task such as flood segmentation, it is particularly important to properly weigh the presence of false-positive pixels incorrectly labeled as “Flood”. A significant number of false positives could suggest a nonexistent flood zone, potentially leading to unwanted triggers. Thus, relying solely on mIoU does not provide sufficient information about false positives, making the F1-Score a important complementary metric.
@@
As shown in Table II, it is noteworthy that all models exhibit relatively similar performance in clouds segmentation, within about 2 percentage points of variation in mIOU between the best and worst, both in FP32 and INT8 quantization, with the exception of UNet + MobileViT-S which suffers greatly from quantization. Furthermore, it is interesting to notice that the SL version of our architecture is only marginally better than the SSL version, suggesting that the design constraint of freezing the backbone may not have a big impact. The differences in performance between the models is even smaller if we look at the F1-Score, indicating again how weel-known SOTA models with higher performance do not have excessive gains in quality metrics compared to more efficient models such as MobileNetV3 and our proposed architecture.
@@
The results on floods segmentation, shown in Table III, show that the higher complexity models are generally superior in this specific task, while the proposed architecture provides better performance than the direct low-complexity alternative (DeepLabV3 + MobileNetV3-L). Indeed, scaling experiments reported in Table VIII suggest that a larger model would improve performance in exchange of speed.
@@
Lastly, the marine littering task addressed a problem of multi-label classification instead of segmentation, providing an analysis of how well the proposed model can address heterogeneous tasks that both need fine and coarse grained features. In this task, we notice that the proposed method is very close to the best FP32 results, and it is the best overall in INT8. It is worth remarking that the MARIDA dataset is very small, leading some highly-complex methods to overfit when finetuned.
@@
Model Tasks Lat. (FP32) Lat. (INT8) Pwr-Norm. Lat. (FP32) Pwr-Norm. Lat. (INT8) Avg Pwr
GhostNetv2 + 3 parallel heads (Ours) SSC 56.77 ms 34.67 ms 48.66 ms 29.72 ms 6.0 W
DeepLabV3 + MobileNetV3-L (DLMN) S 39.52 ms 15.70 ms 32.75 ms 13.01 ms 5.8 W
HRNet18 (HR) S 118.07 ms 47.72 ms 106.26 ms 42.95 ms 6.3 W
UNet + MobileViT Small (mViT) S 117.79 ms 82.94 ms 109.38 ms 77.02 ms 6.5 W
MobileNetV3 Large (MN) C 18.61 ms 9.19 ms 15.59 ms 7.61 ms 5.8 W
ResNet50 (RN50) C 45.98 ms 15.00 ms 42.04 ms 13.71 ms 6.4 W
ViT B-16 (ViT) C 364.79 ms 296.27 ms 343.94 ms 279.34 ms 6.6 W
@@
TABLE VII: Backbone ablation
Backbone Cloud (mIOU) Floods (mIOU) Marine Litter (mF1) Latency
GhostNetV2 (SSL) + 3 parallel heads 82.33 40.32 68.03 56.77 ms
MobileNetV3-L (SSL) + 3 parallel heads 81.45 40.00 64.78 32.51 ms